# arXiv:2602.24141v1[cond-mat.mtrl-sci] 27 Feb 2026

### 27 Feb 2026 v1

@@ II Tight-binding geometry
Carbon nanoscrolls are structurally similar to carbon nanotubes, particularly in the one-round configuration. The key distinction arises from the edge-to-edge hopping, which is interlayer in carbon nanoscrolls and intralayer in carbon nanotubes. The horizontal shift along the axial direction shown in Fig. 1 reconfigures the two atoms previously linked by interlayer hopping to form new interlayer connections with other atoms. This situation is analogous to the geometry of chiral carbon nanotubes, so this geometry is termed a chiral carbon nanoscroll. We therefore adopt the concept of the chiral vector from carbon nanotubes to characterize these structures, defined as follows [26]:
Figure 2: Example of unit cells for one-round and two-round nanoscrolls with chiral vectors $\vec{C}=(-3,6)$, $\vec{C}=(0,6)$, and $\vec{C}=(-1,6)$. Green and purple dots denote lattice sites belonging to different layers. The red line indicates which sites are aligned on top of each other in a scroll configuration. The numbers in the two-round nanoscroll with $\vec{C}=(-3,6)$ correspond to the site labels in Fig. 3. Each unit cell is repeated infinitely along the axial $\hat{x}$ direction, forming a nanoscroll with translational symmetry along its axis.
@@
where $\vec{a}_{1}=(1,0,0)a$ and $\vec{a}_{2}=(\frac{1}{2},\frac{\sqrt{3}}{2},0)a$ are the primitive vectors of graphene. $\vec{\delta}=(0,\frac{1}{\sqrt{3}},0)a$, representing the displacement between A and B sites, is crucial for the ABC-stacking model, as the overlapping atoms in adjacent layers belong to different sublattices. Here, $a$ is the lattice constant, which we set to $1$ for convenience. For simplicity, we denote the chiral vector of a carbon nanoscroll by $(n,m)$. Here $n$ specifies the magnitude of the shift, while $m$ determines the width. All unshifted structures, such as the one shown in Fig. 1(a), satisfy $n=-\frac{m}{2}$.
The chiral vector $\vec{C}$ can be represented on the flattened graphene sheet. Upon rolling the sheet into a multi-round scroll, the two sites connected by the chiral vector will be aligned on top of each other. Based on this definition, Fig. 2 illustrates the unit cells with different chiral vectors for one- and two-round scrolls, all of which exhibit translational symmetry along the $\hat{x}-$direction. Taking the two-round nanoscroll with $\vec{C}=(-3,6)$ in Fig. 2 as an example, the displacement between sites 1 and 12 as well as 13 and 24, indicated by the red line, represents the chiral vector. Fig. 3 illustrates the corresponding rolled-up structure, where sites 1 and 12, as well as 13 and 24, lie above each other when the lattice is rolled up. In addition, all two-round nanoscrolls characterized by $\vec{C}=(n,6)$ exhibit the same cross-sectional geometry in Fig. 3.
@@
III One-round scrolls
We begin our analysis with the band structure of a one-round nanoscroll characterized by the chiral vector $\vec{C}=(-2,12)$. This means that the corresponding unit cell contains 24 sites, and the two edges are shifted with respect to each other by two times the primitive vector $\vec{a}_{1}$ (see the top panels of Fig. 2 for the cases with zero and one shift). Fig. 4(a) shows the band structures for certain values of $\phi$. Dirac cones are observed at $\phi\approx 0.3188$ and $\phi\approx-0.3188$, instead of at $\phi=\pi$, as in Ref. [37]. Since the Hamiltonian exhibits $2\pi$ periodicity in $\phi$, it is natural to introduce a fictitious Brillouin zone (FBZ) spanned by $(k,\phi)$. Fig. 4(b) shows the energy gap at the Fermi level over the FBZ. In this representation, Dirac cones appear as minimum regions (white) enclosed by the yellow circles. The number of Dirac cones agrees with Ref. [18], although their positions in $\phi$ are shifted.
@@ IV Multi-round scrolls
Figure 8: (a) The energy gaps for the two-round nanoscroll with $\vec{C}=(-2,12)$ presented in the FBZ. (b) The band structures under varying magnetic fluxes.
@@ IV.2 Mirror symmetry protection of Dirac cones
$M_{y}$ explains why all Dirac cones are aligned at $\phi=\pi$ in unshifted carbon nanoscrolls. To illustrate this, we take the two-round carbon nanoscroll with $\vec{C}=(-3,6)$ as an example, and adopt a gauge where the phase factor $e^{i\phi}$ is added only to the hoppings between sites “7–8,” “7–18,” and “17–18,” as shown in Fig. 10(a). Fig. 10(b) illustrates that the $M_{y}$ operation, represented as a reflection across the mirror plane, is equivalent to redefining the site indices and reversing the sign of the magnetic flux $\phi$ simultaneously. The matrix $M_{y}$ is anti-diagonal for this choice of site ordering, and further, all its non-zero entries are $1$ under the specific gauge discussed above. One can find a basis that diagonalizes $M_{y}$ into a block-diagonal form $M_{y}^{\prime}$:
@@
where the two blocks correspond to states with different mirror eigenvalues, +1+1 and 1-1. The energy dispersions of $h_{11}$ and $h_{22}$ are shown in Fig. 10(c). The Dirac cones in the figure are formed by the intersection of one band from $h_{11}$ and one from $h_{22}$, indicating that the $M_{y}$ symmetry protects the Dirac cones.
Figure 10: (a) Axial view of the tight-binding model corresponding to a two-round nanoscroll with chiral vector $\vec{C}=(-3,6)$, including the specific gauge for the axial magnetic flux (red hoppings containing the phase factor $e^{i\phi}$). (b) A “flattened” view of the same tight-binding model, indicating the $M_{y}$ mirror operation and the mirror plane. On the right side, we show the axial view of the two lattices, before and after applying $M_{y}$. The transformation of the Hamiltonian by $M_{y}$ is equivalent to redefining the site index starting from the purple site and substituting $\phi$ with $\phi$. Note that the redefinition process corresponds to inverting the stacking sequence of the rounds, assuming that $|1\rangle$ or $|1^{\prime}\rangle$ starts from the inside of the nanoscroll. (c) The energy dispersions of $h_{11}$ and $h_{22}$ in Eq. (12). Each Dirac cone is formed by the intersection of two bands with opposite mirror eigenvalues, demonstrating that they are protected by the $M_{y}$ symmetry.
@@ IV.3 $IT$ symmetry protection of Dirac cones
In shifted carbon nanoscrolls, the mirror symmetry is broken; however, the Dirac cones are preserved. The remaining symmetries are sublattice, spinless time-reversal, and inversion symmetries, as defined in Eqs. (6), (7), and (10). To analyze the effect of sublattice symmetry in protecting the Dirac cones, we break it by considering either an AB potential difference ($V_{a}=-V_{b}=\Delta\neq 0$ ) or the next-nearest-neighbor, intralayer hopping ($t^{\prime}$). The results are shown in Fig. 11. First, we note that an AB sublattice potential difference, which breaks inversion symmetry, introduces a mass term and thus gaps the Dirac cones. In contrast, next-nearest-neighbor hopping $t^{\prime}$ alone preserves the Dirac cones. To better understand the underlying mechanism, we analyze the symmetries defined in Eqs. (7) and (10), which persist even when sublattice symmetry is broken by $t^{\prime}$. By combining these two symmetries, we obtain
@@
where $T={\cal K}$ is the spinless time-reversal operator, equal to complex conjugation. We define a transformed Hamiltonian

### 02 Apr 2026 v2

# Dirac semimetal phases in chiral carbon nanoscrolls

@@ II Tight-binding geometry
Carbon nanoscrolls are structurally similar to carbon nanotubes, particularly in the one-round configuration. The key distinction arises from the edge-to-edge hopping, which is interlayer in carbon nanoscrolls and intralayer in carbon nanotubes. The horizontal shift along the axial direction shown in Fig. 1 reconfigures the two atoms previously linked by interlayer hopping (the red lines in the purple rectangles) to form new interlayer connections with other atoms. This situation is analogous to the chiral carbon nanotubes, so this geometry is termed a chiral carbon nanoscroll. We therefore adopt the chiral vector concept from carbon nanotubes to characterize these structures [26]. On a flattened graphene sheet, the two sites that can be connected by a chiral vector will align on top of each other once the structure is rolled up. The chiral vector in our chiral carbon nanoscroll system is defined as follows:
Figure 2: Example of unit cells for one-round and two-round nanoscrolls with chiral vectors $\vec{C}=(-3,6)$, $(0,6)$, and (1,6)(-1,6). We define the coordinate system such that the $\hat{x}-$axis is oriented along the axial direction of the nanoscroll, the $\hat{y}-$axis corresponds to the transverse direction of the multilayer ribbon, and the $\hat{z}-$axis is normal to the graphene plane. Green and purple dots denote lattice sites belonging to different layers. The red lines highlight pairs of sites within the same layer in the multilayer representation that become vertically aligned in the rolled-up configuration. The numbers in the two-round nanoscroll with $\vec{C}=(-3,6)$ correspond to the site labels in Fig. 3. Each unit cell is repeated infinitely along the axial $\hat{x}$ direction, forming a nanoscroll with translational symmetry.
@@
where $\vec{a}_{1}=(1,0,0)a$ and $\vec{a}_{2}=(\frac{1}{2},\frac{\sqrt{3}}{2},0)a$ are the primitive vectors of graphene. $\vec{\delta}=(0,\frac{1}{\sqrt{3}},0)a$, representing the displacement between A and B sites, is required in the chiral vector definition to satisfy the ABC-stacking condition in our nanoscroll model, as the overlapping atoms in adjacent layers belong to different sublattices. Here, $a$ is the lattice constant, which we set to $1$ for convenience. For simplicity, we denote the chiral vector of a carbon nanoscroll by $(n,m)$. Here $n$ specifies the magnitude of the shift, while $m$ determines the width. All unshifted structures, such as the one shown in Fig. 1(a), satisfy $n=-\frac{m}{2}$.
One can also roll a single-layer graphene into a multi-round scroll. Based on the definition of the chiral vector $\vec{C}$, Fig. 2 illustrates various unit cells for one- and two-round nanoscrolls of large radius within the multilayer representation. All configurations exhibit translational symmetry along the $\hat{x}-$direction, corresponding to the axial direction of the nanoscroll. Here we consider the two-round nanoscroll with $\vec{C}=(-3,6)$ in Fig. 2 as an example. On the flattened multilayer representation, sites connected by $\vec{C}$ within the same layer (e.g., sites 1 and 12, as well as 13 and 24) are mapped onto vertically aligned positions in the rolled-up structure, as illustrated in Fig. 3. We also note that all two-round nanoscrolls characterized by $\vec{C}=(n,6)$ share the same cross-sectional geometry shown in Fig. 3.
@@
III One-round scrolls
We begin our analysis with the band structure of a one-round nanoscroll characterized by the chiral vector $\vec{C}=(-2,12)$. This means that the corresponding unit cell contains 24 sites, and the two edges are shifted with respect to each other by two times the primitive vector $\vec{a}_{1}$ (see the top panels of Fig. 2 for the cases with zero and one shift). Fig. 4(a) shows the band structures for certain values of $\phi$. Dirac cones are observed at $\phi\approx 0.3188$ and $\phi\approx-0.3188$, instead of at $\phi=\pi$, as in Ref. [37]. Since the Hamiltonian exhibits $2\pi$ periodicity in $\phi$, it is natural to introduce a fictitious Brillouin zone spanned by $(k,\phi)$. Fig. 4(b) shows the energy gap at the Fermi level over the FBZ. In this representation, Dirac cones appear as minimum regions (white) enclosed by the yellow circles. The number of Dirac cones agrees with Ref. [18], although their positions in $\phi$ are shifted.
@@ IV Multi-round scrolls
Figure 8: (a) The energy gaps for the two-round nanoscroll with $\vec{C}=(-2,12)$ presented in the FBZ. (b) The band structures under varying magnetic fluxes. The subfigures, presented from left to right, correspond to $\phi=\pi$, 0.3775-0.3775, $1.222$, 1.222-1.222, and $0.3775$, respectively.
@@ IV.2 Mirror symmetry protection of Dirac cones
The mirror symmetry $M_{y}$ explains why all Dirac cones are aligned at $\phi=\pi$ in unshifted carbon nanoscrolls. To illustrate this, we take the two-round carbon nanoscroll with $\vec{C}=(-3,6)$ as an example, and adopt a gauge where the phase factor $e^{i\phi}$ is added only to the hoppings between sites “7–8,” “7–18,” and “17–18,” as shown in Fig. 10(a). Fig. 10(b) illustrates that the $M_{y}$ operation, represented as a reflection across the mirror plane, is equivalent to redefining the site indices and reversing the sign of the magnetic flux $\phi$ simultaneously. The matrix $M_{y}$ is anti-diagonal for this choice of site ordering, and further, all its non-zero entries are $1$ under the specific gauge discussed above. One can find a basis that diagonalizes $M_{y}$ into a block-diagonal form $M_{y}^{\prime}$:
@@
where the two blocks correspond to states with different mirror eigenvalues, +1+1 and 1-1. The energy dispersions of $h_{11}$ and $h_{22}$ are shown in Fig. 10(c). The Dirac cones in the figure are formed by the intersection of one band from $h_{11}$ and one from $h_{22}$, demonstrating that the Dirac cones are protected by the $M_{y}$ symmetry.
Figure 10: (a) Axial view of the tight-binding model corresponding to a two-round nanoscroll with chiral vector $\vec{C}=(-3,6)$, including the specific gauge for the axial magnetic flux (red arrows represent the hoppings with the positive phase factor $e^{i\phi}$). (b) Left: The “flattened” view of the same tight-binding model of the two-round nanoscroll with $\vec{C}=(-3,6)$ in Fig. 2, illustrating the $M_{y}$ mirror operation and the mirror plane. Right: The axial view of the two lattices model, before and after applying $M_{y}$. The transformation of the Hamiltonian by $M_{y}$ is equivalent to redefining the site index starting from the purple site and substituting $\phi$ with $\phi$. Note that the redefinition process corresponds to inverting the stacking sequence of the rounds, assuming that $|1\rangle$ (or $|1^{\prime}\rangle$) and $|24\rangle$ (or $|24^{\prime}\rangle$) denote the innermost and the outermost sites of the nanoscroll, respectively. (c) The energy dispersions of $h_{11}$ and $h_{22}$ in Eq. (12). Each Dirac cone is formed by the intersection of two bands with opposite mirror eigenvalues, demonstrating that they are protected by the $M_{y}$ symmetry.
@@ IV.3 $IT$ symmetry protection of Dirac cones
In shifted carbon nanoscrolls, the mirror symmetry is broken; however, the Dirac cones are preserved. To understand the robustness of the Dirac cones in the absence of mirror symmetry, we examine the remaining symmetries of the system, namely sublattice, spinless time-reversal, and inversion symmetries, as defined in Eqs. (6), (7), and (10). To demonstrate that sublattice symmetry is not essential for the protection of the Dirac cones, we introduce perturbations that break sublattice symmetry, such as an AB potential difference ($V_{a}=-V_{b}=\Delta\neq 0$ ) or next-nearest-neighbor intralayer hopping ($t^{\prime}$). The results are shown in Fig. 11. We find that an AB sublattice potential difference, which breaks inversion symmetry, introduces a mass term and thus gaps the Dirac cones. In contrast, next-nearest-neighbor hopping $t^{\prime}$, which preserves inversion symmetry, does not gap out the Dirac cones. This behavior can be understood in terms of the combined inversion–time-reversal symmetry (IT), under which the Hamiltonian satisfies
@@
where $T={\cal K}$ is the spinless time-reversal operator, equal to complex conjugation. This symmetry protects the Dirac cones even in the absence of sublattice symmetry. To make this symmetry more transparent, we introduce a transformed Hamiltonian